At the initial instant, two particles are observed at different locations moving towards each other with velocities u1 and u2. If they are subjected to constant accelerations a1 and a2 in directions opposite to their initial velocities, they will meet twice. If time interval between these two meetings is Δt, find suitable expression for their initial separation

1
Expert's answer
2021-05-03T10:31:12-0400


Let the initial separation between both the bodies be "x"

Both the bodies will cross each other twice, once during their forward motion and second time during their backward motion.

Taking the body A at rest, so the relative velocity of B with respect to A will be,

"v_{BA}=u_1+u_2"

Relative acceleration of B with respect to A,

"a_{BA}=-(a_1+a_2)" , since acceleration is opposite to the direction of velocity


Time taken by body B to cover the distance "x" will be,

"x=(u_1+u_2)t_1-\dfrac{1}{2}(a_1+a_2)t_1^2\longrightarrow(1)" (this is the first time the two bodies will meet and time instant is "t=t_1" )


Total Time taken by body B to come to rest

"v=u+at_s"

"0=(u_1+u_2)-(a_1+a_2)t_s"

"t_s=\dfrac{u_1+u_2}{a_1+a_2}"

At this time instant the body B has reached the point C from it's initial position

The difference in time interval of the two meetings will be,

"2(t_s-t_1)=\Delta t"

"t_1=t_s-\dfrac{\Delta t}{2}"


Substituting the value of "t_1" and "t_s" in equation (1)

"x=(u_1+u_2)\bigg(\dfrac{u_1+u_2}{a_1+a_2}-\dfrac{\Delta t}{2}\bigg)-\dfrac{1}{2}(a_1+a_2)\bigg(\dfrac{u_1+u_2}{a_1+a_2}-\dfrac{\Delta t}{2}\bigg)^2"

"x=\dfrac{(u_1+u_2)^2}{a_1+a_2}-\dfrac{\Delta t}{2}(u_1+u_2)-\dfrac{1}{2}\dfrac{(u_1+u_2)^2}{a_1+a_2}-\dfrac{(\Delta t)^2}{8}(a_1+a_2)\\\space\space\space\space\space\space\space+\dfrac{\Delta t}{2}(u_1+u_2)"

"x=\dfrac{1}{2}\dfrac{(u_1+u_2)^2}{(a_1+a_2)}-\dfrac{(\Delta t)^2(a_1+a_2)}{8}" (Ans)